volume of O 1So m lithium salfide salutian requiteol te cempletlu reactwith 125.a ml of 0.150 A Selution?
Answer #1

Li2S + Co(NO3)2 -------> CoS + 2LiNO3

No of mole of Co(NO3)2 = (0.150/1000ml)×125ml= 0.01875mole

stoichiometrically, 1mole of Li2S reacting with 1mole of Cu(NO3)2

Therefore,

0.01875mole of Co(NO3)2 require 0.01875 mole of Li2S to completely react

Molarity of Li2S solution = 0.150mol/L

Volume of Li2S solution containing 0.01875mol of Li2S = (1000ml/0.150mol)× 0.01875mol = 125ml